# Pre-trained Universal Medical Image Transformer

### 12 Dec 2023 v1

@@ Conclusion
More Reconstruction Results
We show more visual tokenizer reconstruction results in Figure 5. Our visual tokenizer is able to reconstruct various anatomical structure with diverse imaging modalities and spatial properties.
Classification Experiments
Experimental Settings
Our experimental settings on MedMNIST v2 largely follow the official baselines (Yang et al. 2023), including the regularization for the Adrenal and Vessel datasets. For fair comparisons, we search for optimal learning rate when fine-tuning each model on each dataset. To utilize the pre-trained weights of models, we adopt the following adjustments. First, for each model, each input image is interpolated into the size that is compatible with the input size used by the model during pre-training. Second, the 3D images of MedMNIST v2 have only one channel. For each model that is pre-trained on images with three channels (such as RGB images), we duplicate each input image along the channel dimension into three channels. The detailed recipe is provided in Table 5.
Segmentation Experiments
Experimental Settings
@@
where $d(s,t)$ is the Euclidean distance between points $s$ and $t$. More specific details of the evaluation metrics can be found in the official documentation of the CHAOS challenge [^2] https://github.com/emrekavur/CHAOS-evaluation/blob/master/CHAOS˙Metrics˙and˙Evaluation˙Method.pdf. For results on BTCV validation set, we use the MONAI’s implementation for the metrics to evaluate them. For results on CHAOS test set, we submit them to the challenge [^3] https://chaos.grand-challenge.org and report the evaluation results returned by the website.

### 24 Jan 2024 v2

@@ Appendix C Pre-training Data and Pre-processing
Appendix D More Reconstruction Results
We show more reconstruction results by SPAD-VT in Figure 6. SPAD-VT can reconstruct various anatomical structure with diverse imaging modalities and spatial properties.
Appendix E Classification Experiments
Our experimental settings on MedMNIST v2 largely follow the official baselines Yang et al. [2023], including the regularization for the Adrenal and Vessel datasets. For fair comparisons, we search for optimal learning rate when fine-tuning each model on each dataset. To utilize the pre-trained weights of models, we adopt the following adjustments. First, for each model, each input image is interpolated into the size that is compatible with the input size used by the model during pre-training. Second, the 3D images of MedMNIST v2 have only one channel. For each model that is pre-trained on images with three channels (such as RGB images), we duplicate each input image along the channel dimension into three channels. The detailed recipe is provided in Table 6.
Appendix F Segmentation Experiments
F.1 Experimental Settings
@@ F.2 Evaluation Metrics
where $d(s,t)$ is the Euclidean distance between points $s$ and $t$. More specific details of the evaluation metrics can be found in the official documentation of the CHAOS challenge [^2] https://github.com/emrekavur/CHAOS-evaluation/blob/master/CHAOS_Metrics_and_Evaluation_Method.pdf. For results on BTCV validation set, we use the MONAI’s implementation for the metrics to evaluate them. For results on CHAOS test set, we submit them to the challenge [^3] https://chaos.grand-challenge.org and report the evaluation results returned by the website.